Please help me solve for x. I keep getting stuck.
f(x)= x^4+x^2-12.

Answer by edjones(8007)

f(x)= x^4+x^2-12
=(x^2+4)(x^2-3)
Zeroes for x:
x^2=3
x=+-sqrt(3)
.
x^2=-4
x=+-sqrt(-4)
x=+-2i imaginary numbers.
.
Ed